A Bibliography of Supercomputing ’2004

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

12 April 2006 Version 1.01

Abstract

This bibliography records articles presented at the Supercomputing ’2004 conference.

Title word cross-reference

4 [WCSK04]. 5 [FC04].

-D [FC04]. -Dimensional [WCSK04].

10 [BCC^{+}04]. 15.2 [KKF^{+}04].

2 [MK04]. 2004 [ACM04].

Abundant [Won04]. ActiveScale [TGZ^{+}04a]. Adaptive [DA04]. Advanced [USB04]. Aerospace [MTA^{+}04]. Algorithms [ORR04].

Alternative [CKE^{+}04]. Analyses [ABKP04]. Analysis [USB04]. Application [HCB04]. Application-Aware [LOT04]. Application-Level [SBF^{+}04]. Applications [PDT^{+}04]. Architecture [SCY04]. Architecture-Aware [PTCV04]. Architectures [PDT^{+}04]. Assessing [dLR04]. Automated [CH04]. Automatic [GAW04]. Automatically [TH04]. Aware [PTCV04].

Bandwidth [TGZ^{+}04a]. Based [FPP04]. Be [DeB04]. Big [LOT04]. Billion [ABKP04]. Billions [TO04]. Biological [Won04]. Blade [PPH^{+}04]. BlueGene [DHJ^{+}04]. BlueGene/L [DHJ^{+}04]. Boltzmann [PDT^{+}04]. Bridging [ACM04].

Building [LVP04]. Bundle [ORR04].

Cache [BH04]. Caches [LOT04]. Caching [ORR04]. Calculations [BCC^{+}04]. Case [EAK04]. Centric [YF04]. CFD [MTA^{+}04]. Checkpoint [SBF^{+}04]. Checkpoint-Recovery [SBF^{+}04]. Cluster [Bur04, TGZ^{+}04a]. Cluster-Based [SCY04]. Clusters [CKE^{+}04]. Collaborative [GAW04]. Communication [CG04]. communities [ACM04]. Compiler [PE04]. Complex [PTCV04]. Compression [KBS04]. Computation [DO04]. Computational [TO04]. Computations [OCC^{+}04]. Computer [BCC^{+}04]. Computing [KGZ^{+}04]. Concept [PPH^{+}04]. Conference [ACM04]. Connected [DHL^{+}04]. Context [Hol04]. Control [SH04]. Coordination [KNT^{+}04]. Coscheduling [CKE^{+}04]. Counters [TH04]. Coupled [BCC^{+}04].

D [FC04]. Data [ORR04]. Data-Grids [ORR04]. Data-Intensive [TGZ^{+}04b]. Database [TO04]. Datasets [WCSK04]. Degrees [ABKP04]. Delivering [TGZ^{+}04a]. Departmental [PPH^{+}04]. Design [USB04]. Desktop [KCC04b]. Dimensional [WCSK04]. Discovery [Ree04]. Disk [Bur04]. Disk-Resident [WCSK04]. Distributed [CG04]. Distribution [GAW04]. Dynamic [WPBM04]. Dynamics [BCF^{+}04].

Early [MTA^{+}04]. Earth [KKF^{+}04]. Earthquake [YMW04]. Economy [Aha04].

Efficient [GHW^{+}04]. Element [ABKP04]. Elements [TO04]. Enabled [GAW04]. Enhancement [BJZH04]. Enterprise [KCC04b]. Environment [GAW04]. Environments [KGZ^{+}04]. Era [Won04]. Evaluating [CG04]. Evaluation [PDT^{+}04]. Execution [KGZ^{+}04]. Existing [EAK04]. Experience [MTA^{+}04]. Experiences [GHG04].

Fastpath [Bur04]. Fat [KNT^{+}04].

Fault [dLR04]. Fault-Tolerant [DHL^{+}04]. File [WPBM04]. File-Bundle [ORR04]. Finite [ABKP04]. Flexible [FPP04]. Flow [SH04]. Fluid [SH04]. Framework [RR04]. Freedom [ABKP04]. Fujitsu [MTA^{+}04]. Functionally [SCY04].

Generatinn [TO04]. Geodynamo [KKF^{+}04]. Geometric [RR04]. GPU [FQKYS04]. Grid [GAW04]. Grids [ORR04]. GYRO [FC04]. Gyrokinetic [FC04]. Gyrokinetic-Maxwell [FC04].

Half [ABKP04]. Haralick [WCSK04]. Hardware [HKR04]. Harness [SOE^{+}04]. Hash [CCF04]. Hexahedral [TO04]. High [HCB04]. HPC2500 [MTA^{+}04].

IEEE [ACM04]. IEEE/ACM [ACM04]. Image [WCSK04]. Implementation [GHG04]. Implicit [ABKP04]. Improve [CH04]. Improving [DO04]. Inca [SOE^{+}04]. InfiniBand [LVP04]. Information [CH04]. Intel [BH04]. Intellectual [Ree04]. Intensive [TGZ^{+}04b]. Inter [KNT^{+}04]. Inter-Layer [KNT^{+}04]. Internet [DHL^{+}04]. Itanium [BH04].

Java [HCB04]. JAXA [MTA^{+}04].

Kosha [BJZH04].

L [DHJ^{+}04]. Lambda [YF04]. Lambda-Grid [YF04]. Language [DA04]. Large [PDT^{+}04]. Large-Scale [PDT^{+}04]. Lattice [PDT^{+}04]. Law [DeB04]. Layer [KNT^{+}04]. Level [RR04]. Lever [Ree04]. Line [BCF^{+}04]. Linux [MK04]. Locality [DO04]. Location [CCF04]. Logistics [Swa04]. Long [KNT^{+}04].

Machine [KGZ^{+}04]. Management [WPBM04]. Managing [KGZ^{+}04]. Maxwell [FC04]. Measurement [BH04]. Mechanics [ABKP04]. Memory [TH04]. Merrimac [EAG^{+}04]. Meshes [TO04]. Messages [KBS04]. Messaging [FPP04]. Metadata [WPBM04]. Middleware [HCB04]. Modeling [USB04]. Modern [OCC^{+}04]. Modernizing [EAK04]. Molecular [EAG^{+}04]. Moore [DeB04]. MPI [MK04]. MPI-2 [MK04]. MPI-Level [LVP04]. Multi [RR04]. Multi-Level [RR04]. Multidisciplinary [Ree04]. Multirail [LVP04].

Network [SCY04, BJZH04]. Networking [ACM04]. Networks [KNT^{+}04]. Nodes [DHL^{+}04]. November [ACM04].

Online [LC04]. Opportunities [Won04]. Optimal [ORR04]. Optimization [PTCV04]. Optimizations [Bur04]. Organizing [TGZ^{+}04b].

PA [ACM04]. Panasas [TGZ^{+}04a]. ParaDiS [BCF^{+}04]. Parallel [OCC^{+}04]. Parallelization [GHW^{+}04].

Peer [BJZH04]. Peer-to-Peer [BJZH04]. ACM [ACM04]. Performance [HCB04]. Petabyte [WPBM04]. Petabyte-Scale [WPBM04]. PIM [USB04]. Pipe [KNT^{+}04]. Pipeline [YMW04]. Pittsburgh [ACM04]. Pools [HKR04]. Post [GHW^{+}04]. Post-Processing [GHW^{+}04]. Potential [DO04]. Pragmatic [PTCV04]. Predicting [CG04]. PRIMEPOWER [MTA^{+}04]. Prior [CH04]. Proceedings [ACM04]. Process [PTCV04]. Processing [GHW^{+}04]. Processor [BH04]. Processors [PTCV04]. Program [PTCV04]. Programming [RR04]. Programs [SBF^{+}04]. Proiectiles [SH04]. Protocol [GHG04]. Providing [KGZ^{+}04].

QCDOC [BCC^{+}04].

Rapid [KCC04b]. Rating [PE04]. Real [HCB04]. Realistic [LC04]. Recovery [Bur04]. Regrouping [DO04]. Rendering [GAW04]. Replica [CCF04]. Reporting [SOE^{+}04]. Resident [WCSK04]. Resource [KCC04b]. Resources [KCC04a]. Results [EAG^{+}04]. Routed [YF04]. RPC [DHL^{+}04]. RPC-V [DHL^{+}04]. Runs [CH04]. Runtime [KBS04].

Scalability [DHJ^{+}04]. Scalable [BCF^{+}04]. Scale [WPBM04]. Scheme [SBF^{+}04]. Scientific [OCC^{+}04]. Self [TGZ^{+}04b]. Self-Organizing [TGZ^{+}04b]. Sensitivity [dLR04]. Service [CCF04]. Services [FPP04]. Shared [Bur04]. Shared-Disk [Bur04]. Simulation [LC04]. Simulations [SH04]. Simulator [KKF^{+}04]. Small [LOT04]. SOAP [FPP04]. SOAP-Based [FPP04]. Software [EAK04]. Solid [ABKP04]. Solver [FC04]. Storage [TGZ^{+}04b]. Streams [KNT^{+}04]. Study [EAK04]. Sufficient [DeB04].

Supercomputer [ACG^{+}04]. Supercomputing [PDT^{+}04]. Support [MK04]. Supporting [SCY04]. Symmetric [SCY04]. Synchronization [HKR04]. Synthesis [KCC04a]. System [TO04, BJZH04]. Systematic [PTCV04]. Systems [WPBM04].

Table [CCF04]. Task [HKR04]. TCP [KNT^{+}04]. Teraflops [BCC^{+}04]. Terascale [YMW04]. Test [SOE^{+}04]. Texture [WCSK04]. TFlops [KKF^{+}04]. Three [PDT^{+}04]. Throughput [Swa04]. Tightly [BCC^{+}04]. Tightly-Coupled [BCC^{+}04]. Tiling [RR04]. Tolerant [DHL^{+}04]. Tool [MK04]. Tradeoffs [USB04]. Transport [GHG04]. Tuning [PE04]. Turnaround [KCC04b].

Ultrascalable [ABKP04]. Unlocking [ACG^{+}04]. Unstructured [TO04]. User [YF04]. User-Centric [YF04]. Using [CH04].

V [DHL^{+}04]. Vector [OCC^{+}04]. Vehicle [PPH^{+}04]. Viable [CKE^{+}04]. VIRACOCHA [GHW^{+}04]. Virtual [KGZ^{+}04]. Visualization [GAW04]. VMPlants [KGZ^{+}04]. Volatile [DHL^{+}04]. VTOL [PPH^{+}04].

Wavelength [YF04]. Wavelength-Routed [YF04]. Will [DeB04]. Wins [LOT04]. Workloads [GAW04].

References

Adams:2004:UIF

[ABKP04] Mark F. Adams, Harun H.

Bayraktar, Tony M. Keaveny, and Panayiotis Papadopoulos. Ultrascalable implicit finite element analyses in solid mechanics with over a half a billion degrees of freedom. In ACM [ACM04], page 34. ISBN 0-7695-2153-3. LCCN ????

Almasi:2004:UPB

[ACG^{+}04] George Almasi, Siddhartha Chatterjee, Alan Gara, John Gunnels, Manish Gupta, Amy Henning, Jose E. Moreira, and Bob Walkup. Unlocking the performance of the BlueGene/L supercomputer. In ACM [ACM04], page 57. ISBN 0-7695-2153-3. LCCN ????

ACM:2004:SHP

[ACM04] ACM, editor. SC 2004: High Performance Computing, Networking and Storage: Bridging communities: Proceedings of the IEEE/ACM Supercomputing 2004 Conference, Pittsburgh, PA, November 6–12, 2004. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2004. ISBN 0-7695-2153-3. LCCN ????

Ahalt:2004:THP

[Aha04] Stan Ahalt. Toward a high performance computing economy. In ACM [ACM04], page 61. ISBN 0-7695-2153-3. LCCN ????

Boyle:2004:QTC

[BCC^{+}04] P. A. Boyle, Dong Chen, Norman H. Christ, Mike Clark, Saul Cohen, Zhihua Dong, Alan Gara, Balint Joo, Chulwoo Jung, Ludmila Levkova, Xiaodong Liao, Guofeng Liu, Robert D. Mawhinney, Shigemi Ohta, Konstantin Petrov, Tilo Wettig, Azusa Yamaguchi, and Calin Cristian. QC-DOC: A 10 Teraflops computer for tightly-coupled calculations. In ACM [ACM04], page 40. ISBN 0-7695-2153-3. LCCN ????

Bulatov:2004:SLD

[BCF^{+}04] Vasily Bulatov, Wei Cai, Jeff Fier, Masato Hiratani, Gregg Hommes, Tim Pierce, Meijie Tang, Moono Rhee, Kim Yates, and Tom Arsenlis. Scalable line dynamics in ParaDiS. In ACM [ACM04], page 19. ISBN 0-7695-2153-3. LCCN ????

Buck:2004:DCC

[BH04] Bryan R. Buck and Jeffrey K.

Hollingsworth. Data centric cache measurement on the Intel Itanium 2 processor. In ACM [ACM04], page 58. ISBN 0-7695-2153-3. LCCN ????

Butt:2004:KPP

[BJZH04] Ali Raza Butt, Troy A. Johnson, Yili Zheng, and Y. Charlie Hu. Kosha: A peer-to-peer enhancement for the Network File System. In ACM [ACM04], page 51. ISBN 0-7695-2153-3. LCCN ????

Burns:2004:FOC

[Bur04] Randal Burns. Fastpath optimizations for cluster recovery in shared-disk systems. In ACM [ACM04], page 5. ISBN 0-7695-2153-3. LCCN ????

Cai:2004:PPR

[CCF04] Min Cai, Ann Chervenak, and Martin Frank. A peer-to-peer replica location service based on a distributed hash table. In ACM [ACM04], page 56. ISBN 0-7695-2153-3. LCCN ????

Cameron:2004:PED

[CG04] Kirk W. Cameron and Rong Ge. Predicting and evaluating distributed communication performance. In ACM [ACM04], page 43. ISBN 0-7695-2153-3. LCCN ????

Chung:2004:UIP

[CH04] I-Hsin Chung and Jeffrey K. Hollingsworth. Using information from prior runs to improve automated tuning systems. In ACM [ACM04], page 30. ISBN 0-7695-2153-3. LCCN ????

Choi:2004:CCI

[CKE^{+}04] Gyu Sang Choi, Jin-Ha Kim, Deniz Ersoz, Andy B. Yoo, and Chita R. Das. Coscheduling in clusters: Is it a viable alternative? In ACM [ACM04], page 16. ISBN 0-7695-2153-3. LCCN ????

Du:2004:LCS

[DA04] Wei Du and Gagan Agrawal. Language and compiler support for adaptive applications. In ACM [ACM04], page 29. ISBN 0-7695-2153-3. LCCN ????

DeBenedictis:2004:WML

[DeB04]  Erik P. DeBenedictis.  Will

Moore’s Law be sufficient? In ACM [ACM04], page 45. ISBN 0-7695-2153-3. LCCN ????

Davis:2004:PSA

[DHJ^{+}04] Kei Davis, Adolfy Hoisie, Greg Johnson, Darren J. Kerbyson, Mike Lang, Scott Pakin, and Fabrizio Petrini. A performance and scalability analysis of the Blue-Gene/L architecture. In ACM [ACM04], page 41. ISBN 0-7695-2153-3. LCCN ????

Djilali:2004:RVT

[DHL^{+}04] Samir Djilali, Thomas Herault, Oleg Lodygensky, Tangui Morlier, Gilles Fedak, and Franck Cappello. RPC-V: Toward faulttolerant RPC for Internet connected desktop grids with volatile nodes. In ACM [ACM04], page 39. ISBN 0-7695-2153-3. LCCN ????

Lu:2004:AFS

[dLR04] Charng da Lu and Daniel A.

Reed. Assessing fault sensitivity in MPI applications. In ACM [ACM04], page 37. ISBN 0-7695-2153-3. LCCN ????

Ding:2004:PCR

[DO04] Chen Ding and Maksim Orlovich.

The potential of computation regrouping for improving locality. In ACM [ACM04], page 13. ISBN 0-7695-2153-3. LCCN ????

Erez:2004:APR

[EAG^{+}04] Mattan Erez, Jung Ho Ahn, Ankit Garg, William J. Dally, and Eric Darve. Analysis and performance results of a molecular modeling application on Merrimac. In ACM [ACM04], page 42. ISBN 0-7695-2153-3. LCCN ????

Everaars:2004:MES

[EAK04] C. T. H. Everaars, F. Arbab, and B. Koren. Modernizing existing software: A case study. In ACM [ACM04], page 3. ISBN 0-7695-2153-3. LCCN ????

Fahey:2004:GDG

[FC04] Mark R. Fahey and Jeff Candy. GYRO: A 5-D gyrokinetic-Maxwell solver. In ACM [ACM04], page 26. ISBN 0-7695-2153-3. LCCN ????

Fox:2004:TFM

[FPP04] Geoffrey Fox, Shrideep Pallickara, and Savas Parastatidis. Toward flexible messaging for SOAP-based services. In ACM [ACM04], page 8. ISBN 0-7695-2153-3. LCCN ????

Fan:2004:GCH

[FQKYS04] Zhe Fan, Feng Qiu, Arie Kaufman, and Suzanne Yoakum-Stover. GPU cluster for high performance computing. In ACM [ACM04], page 47. ISBN 0-7695-2153-3. LCCN ????

Grimstead:2004:ADR

[GAW04] Ian J. Grimstead, Nick J. Avis, and David W. Walker. Automatic distribution of rendering workloads in a Grid enabled collaborative visualization environment. In ACM [ACM04], page 1. ISBN 0-7695-2153-3. LCCN ????

Gu:2004:EDI

[GHG04] Yunhong Gu, Xinwei Hong, and Robert L. Grossman. Experiences in design and implementation of a high performance transport protocol. In ACM [ACM04], page 22. ISBN 0-7695-2153-3. LCCN ????

Gerndt:2004:VEP

[GHW^{+}04] Andreas Gerndt, Bernd Hentschel, Marc Wolter, Torsten Kuhlen, and Christian Bischof. VIRA-COCHA: An efficient parallelization framework for large-scale CFD post-processing in virtual environments. In ACM [ACM04], page 50. ISBN 0-7695-2153-3. LCCN ????

Huet:2004:HPJ

[HCB04] Fabrice Huet, Denis Caromel, and Henri E. Bal. A high performance Java middleware with a real application. In ACM [ACM04], page 2. ISBN 0-7695-2153-3. LCCN ????

Hoffmann:2004:PET

[HKR04] Ralf Hoffmann, Matthias Korch, and Thomas Rauber. Performance evaluation of task pools based on hardware synchronization. In ACM [ACM04], page 44. ISBN 0-7695-2153-3. LCCN ????

Holland:2004:HPC

[Hol04] Charles J. Holland. High performance computing in context. In ACM [ACM04], page 60. ISBN 0-7695-2153-3. LCCN ????

Ke:2004:RCM

[KBS04] Jian Ke, Martin Burtscher, and Evan Speight. Runtime compression of MPI messages to improve the performance and scalability of parallel applications. In ACM [ACM04], page 59. ISBN 0-7695-2153-3. LCCN ????

Kee:2004:RMS

[KCC04a] Yang-Suk Kee, Henri Casanova, and Andrew A. Chien. Realistic modeling and synthesis of resources for computational grids. In ACM [ACM04], page 54. ISBN 0-7695-2153-3. LCCN ????

Kondo:2004:RMR

[KCC04b] Derrick Kondo, Andrew A. Chien, and Henri Casanova. Resource management for rapid application turnaround on enterprise desktop grids. In ACM [ACM04], page 17. ISBN 0-7695-2153-3. LCCN ????

Krsul:2004:VPM

[KGZ^{+}04] Ivan Krsul, Arijit Ganguly, Jian Zhang, Jose A. B. Fortes, and Renato J. Figueiredo. VMPlants: Providing and managing virtual machine execution environments for Grid computing. In ACM [ACM04], page 7. ISBN 0-7695-2153-3. LCCN ????

Kageyama:2004:TSG

[KKF^{+}04] Akira Kageyama, Masanori Kameyama, Satoru Fujihara, Masaki Yoshida, Mamoru Hyodo, and Yoshinori Tsuda. A 15.2 TFlops simulation of Geodynamo on the Earth Simulator. In ACM [ACM04], page 35. ISBN 0-7695-2153-3. LCCN ????

Kamezawa:2004:ILC

[KNT^{+}04] Hiroyuki Kamezawa, Makoto Nakamura, Junji Tamatsukuri, Nao Aoshima, Mary Inaba, and Kei Hiraki. Inter-layer coordination for parallel TCP streams on long fat pipe networks. In ACM [ACM04], page 24. ISBN 0-7695-2153-3. LCCN ????

Liu:2004:RLS

[LC04] Xin Liu and Andrew A. Chien.

Realistic large-scale online network simulation. In ACM [ACM04], page 31. ISBN 0-7695-2153-3. LCCN ????

Lopez:2004:BWS

[LOT04] Julio C. Lopez, David R.

O’Hallaron, and Tiankai Tu. Big wins with small applicationaware caches. In ACM [ACM04], page 20. ISBN 0-7695-2153-3. LCCN ????

Liu:2004:BMI

[LVP04] Jiuxing Liu, Abhinav Vishnu, and Dhabaleswar K. Panda. Building multirail InfiniBand clusters: MPI-level design and performance evaluation. In ACM [ACM04], page 33. ISBN 0-7695-2153-3. LCCN ????

Mohror:2004:PTS

[MK04] Kathryn Mohror and Karen L.

Karavanic. Performance tool support for MPI-2 on Linux. In ACM [ACM04], page 28. ISBN 0-7695-2153-3. LCCN ????

Matsuo:2004:EEA

[MTA^{+}04] Yuichi Matsuo, Masako Tsuchiya, Masaki Aoki, Naoki Sueyasu, Tomohide Inari, and Katsumi Yazawa. Early experience with aerospace CFD at JAXA on the Fujitsu PRIMEPOWER HPC2500. In ACM [ACM04], page 11. ISBN 0-7695-2153-3. LCCN ????

Oliker:2004:SCM

[OCC^{+}04] Leonid Oliker, Andrew Canning, Jonathan Carter, John Shalf, and Stephane Ethier. Scientific computations on modern parallel vector systems. In ACM [ACM04], page 10. ISBN 0-7695-2153-3. LCCN ????

Otoo:2004:OFB

[ORR04] Ekow Otoo, Doron Rotem, and Alexandru Romosan. Optimal file-bundle caching algorithms for data-grids. In ACM [ACM04], page 6. ISBN 0-7695-2153-3. LCCN ????

Pohl:2004:PEP

[PDT^{+}04] Thomas Pohl, Frank Deserno, Nils Thurey, Ulrich Rude, Peter Lammers, Gerhard Wellein, and Thomas Zeiser. Performance evaluation of parallel large-scale lattice Boltzmann applications on three supercomputing architectures. In ACM [ACM04], page 21. ISBN 0-7695-2153-3. LCCN ????

Pan:2004:RCO

[PE04] Zhelong Pan and Rudolf Eigenmann. Rating compiler optimizations for automatic performance tuning. In ACM [ACM04], page 14. ISBN 0-7695-2153-3. LCCN ????

Park:2004:OBS

[PPH^{+}04] Jin Woo Park, Si Hyoung Park, In Seong Hwang, Ji Joong Moon, Youngha Yoon, and Seung Jo Kim. Optimal blade system design of a new concept VTOL vehicle using the departmental computing Grid system. In ACM [ACM04], page 36. ISBN 0-7695-2153-3. LCCN ????

Parello:2004:TSP

[PTCV04] David Parello, Olivier Temam, Albert Cohen, and Jean-Marie Verdun. Towards a systematic, pragmatic and architectureaware program optimization process for complex processors. In ACM [ACM04], page 15. ISBN 0-7695-2153-3. LCCN ????

Reed:2004:CIL

[Ree04] Daniel A. Reed. Computing — an intellectual lever for multidisciplinary discovery. In ACM [ACM04], page 63. ISBN 0-7695-2153-3. LCCN ????

Renganarayana:2004:GPF

[RR04] Lakshminarayanan Renganarayana and Sanjay Rajopadhye. A geometric programming framework for optimal multi-level tiling. In ACM [ACM04], page 18. ISBN 0-7695-2153-3. LCCN ????

Schulz:2004:IES

[SBF^{+}04] Martin Schulz, Greg Bronevetsky, Rohit Fernandes, Daniel Marques, Keshav Pingali, and Paul Stodghill. Implementation and evaluation of a scalable application-level checkpointrecovery scheme for MPI programs. In ACM [ACM04], page 38. ISBN 0-7695-2153-3. LCCN ????

Shen:2004:SCB

[SCY04] Kai Shen, Lingkun Chu, and Tao Yang. Supporting cluster-based network services on functionally symmetric software architecture. In ACM [ACM04], page 9. ISBN 0-7695-2153-3. LCCN ????

Sahu:2004:ACF

[SH04] Jubaraj Sahu and Karen R. Heavey. Advanced computational fluid dynamics simulations of proiectiles with flow control. In ACM [ACM04], page 27. ISBN 0-7695-2153-3. LCCN ????

Smallen:2004:ITH

[SOE^{+}04] Shava Smallen, Catherine Olschanowsky, Kate Ericson, Pete Beckman, and Jennifer M. Schopf. The Inca test harness and reporting framework. In ACM [ACM04], page 55. ISBN 0-7695-2153-3. LCCN ????

Swany:2004:ITG

[Swa04]  Martin Swany.  Improving

throughput for Grid applications with network logistics. In ACM [ACM04], page 23. ISBN 0-7695-2153-3. LCCN ????

Tang:2004:PAS

[TGZ^{+}04a] Hong Tang, Aziz Gulbeden, Jingyu Zhou, William Strathearn, Tao Yang, and Lingkun Chu. The Panasas ActiveScale Storage Cluster — delivering scalable high bandwidth storage. In ACM [ACM04], page 53. ISBN 0-7695-2153-3. LCCN ????

Tang:2004:SOS

[TGZ^{+}04b] Hong Tang, Aziz Gulbeden, Jingyu Zhou, William Strathearn, Tao Yang, and Lingkun Chu. A self-organizing storage cluster for parallel data-intensive applications. In ACM [ACM04], page 52. ISBN 0-7695-2153-3. LCCN ????

Tikir:2004:UHC

[TH04] Mustafa M. Tikir and Jeffrey K.

Hollingsworth. Using hardware counters to automatically improve memory performance. In ACM [ACM04], page 46. ISBN 0-7695-2153-3. LCCN ????

Tu:2004:CDS

[TO04] Tiankai Tu and David R.

O’Hallaron. A computational database system for generatinn unstructured hexahedral meshes with billions of elements. In ACM [ACM04], page 25. ISBN 0-7695-2153-3. LCCN ????

Upchurch:2004:AMA

[USB04] Ed Upchurch, Thomas Sterling, and Jay Brockman. Analysis and modeling of advanced PIM architecture design tradeoffs. In ACM [ACM04], page 12. ISBN 0-7695-2153-3. LCCN ????

Woods:2004:PID

[WCSK04] Brent Woods, Bradley Clymer, Joel Saltz, and Tahsin Kurc. A parallel implementation of 4-dimensional Haralick texture analysis for disk-resident image datasets. In ACM [ACM04], page 48. ISBN 0-7695-2153-3. LCCN ????

Wong:2004:COE

[Won04] Gane Ka-Shu Wong. Computing opportunities in the era of abundant biological data. In ACM [ACM04], page 62. ISBN 0-7695-2153-3. LCCN ????

Weil:2004:DMM

[WPBM04] Sage A. Weil, Kristal T. Pollack, Scott A. Brandt, and Ethan L. Miller. Dynamic metadata management for petabyte-scale file systems. In ACM [ACM04], page 4. ISBN 0-7695-2153-3. LCCN ????

Yu:2004:CUC

[YF04] Oliver T. Yu and Thomas A. De Fanti. Collaborative user-centric lambda-Grid over wavelengthrouted network. In ACM [ACM04], page 32. ISBN 0-7695-2153-3. LCCN ????

Yu:2004:PVP

[YMW04] Hongfeng Yu, Kwan-Liu Ma, and Joel Welling. A parallel visualization pipeline for terascale earthquake simulations. In ACM [ACM04], page 49. ISBN 0-7695-2153-3. LCCN ????